OpenStudy (anonymous):

Wilma drove at an average speed of 60mph from her home in City A to visit her sister in City B. She stayed in City B 25 hours, and on the trip back averaged 40mph. She returned home 42hours after leaving. We have to determine how many miles x is it from City A to City B. 1. In terms of x, express Wilma's driving time from A to B It is ____________ 2. In terms of x, her return driving time is ______________ 3. Her TOTAL driving time is: _____________ (trip time minus visiting time) So answer from 1. PLUS 2. set equal to 3. is the equation: _______________________ = ___________________.

OpenStudy (anonymous):

This is the problem...

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

I'm not sure if I enter these in wrong or what.. but are you able to tell me which number goes to which blank?

OpenStudy (anonymous):

#1 avg speed=60mph; distance=x, then time=x/60 #2 time=x/40 #3 total driving time=42-25 x can be obtained by solving: x/60+x/40=42-25

OpenStudy (anonymous):

I've entered what annand have and it's correct.. but I'm trying to solve for the last problem now. which is solving for x

OpenStudy (anonymous):

exactly, suzi... but isn't 42=(travel time from a to b)+ staying time+(ravel time from b to a)

OpenStudy (anonymous):

so to solve X, (x/60) + (x/40) = 17 I would have to multiply 60 by each side, correct? and solve?

OpenStudy (anonymous):

better multiply both sides with 120... its a multiple of both 40 & 60...

OpenStudy (anonymous):

Are you able to show me the steps?

OpenStudy (anonymous):

try your hand... if its difficult, i'll help...

OpenStudy (anonymous):

17*120 = 2040.. right? then does that make the x/60 + x/40 = to x? since it all cancels out?

OpenStudy (anonymous):

after multiplying with 120: 2x+3x=17*120 5x=17*120 x=17*120/5

OpenStudy (anonymous):

you don't need to find 17*120 outrightly... 120/5=24

OpenStudy (anonymous):

so how did you get 3 and 2?

OpenStudy (anonymous):

2x + 3x

OpenStudy (anonymous):

Thanks a lot annand! <3

OpenStudy (anonymous):

\[120*\left( x/60+x/40 \right)=(120*x/60)+(120*x/40)=2x+3x\]

OpenStudy (anonymous):

:-)